### v1

@@ 1.1 Angular-Momentum Scaling in a Turbulent Flow
where the subscripts ”1” and ”2” indicate quantities at two points $\bm{r}_{1}$ and $\bm{r}_{2}$, respectively. Under the assumption of statistical homogeneity and isotropy, it is straightforward to show that $\langle\rho_{1}\rho_{2}v_{1i}v_{2j}\rangle=B^{\rho,v}_{ij}(\bm{s})-\frac{1}{2}B_{\rho}S^{\rm dw}_{ij}(\bm{s})$, where the density correlation function $B_{\rho}(\bm{s})\equiv\langle\rho_{1}\rho_{2}\rangle$, the mixed correlation function $B^{\rho,v}_{ij}(\bm{s})\equiv\langle\rho_{1}\rho_{2}v_{1i}v_{1j}\rangle$ and the density-weighted structure function $S^{\rm dw}_{ij}(\bm{s})\equiv\langle\rho_{1}\rho_{2}(v_{2i}-v_{1i})(v_{2j}-v_{1j})\rangle/B_{\rho}$, depend on the separation, $\bm{s}=\bm{r}_{2}-\bm{r}_{1}$. If the density and velocity fields are assumed to be independent, we have $B^{\rho,v}_{ij}=\langle\rho_{1}\rho_{2}\rangle\langle v_{1i}v_{1j}\rangle=B_{\rho}v^{\prime 2}\delta_{ij}$ with $v^{\prime}$ the 1-dimensional rms velocity, suggesting that the longitudinal and transverse correlation functions $B^{\rho,v}_{ll}=B^{\rho,v}_{nn}=B_{\rho}v^{\prime 2}$ (see Figure 4 and discussions below). The above equation can then be rewritten as,
@@
Figure 4: Left: Longitudinal (red squares) and transverse (blue circles) components of the mixed correlation function $B^{\rho,v}_{ij}=\langle\rho_{1}\rho_{2}v_{1i}v_{2j}\rangle$. Under the assumption of independence between $\rho$ and ${\bm{v}}$, both are expected to equal $v^{\prime 2}B_{\rho}$ (black circles). The density correlation function, $B_{\rho}$, exhibits a power-law scaling with an exponent of 0.61 in the inertial range. Right: Longitudinal (circles) and transverse (squares) components the velocity structure tensor with (red) and without (blue) density weighting. The structure functions show similar scalings, but the amplitudes of the density-weighted ones are slightly smaller.
The two contributions in equation (13) can be intuitively understood as follows. The first term which arises mainly due to density fluctuations represents the offset of the mass center from the geometric center of sphere (see below), while the second term, which depends on the velocity structure function, originates from the ”imbalance” of the turbulent velocity on the opposite sides of the geometric center, leading to a ”residual” angular momentum. For convenience, we denote the two terms as $\langle\bm{J}^{2}\rangle_{1}$ and $\langle\bm{J}^{2}\rangle_{2}$, respectively. As discussed below, the first term provides the dominant contribution in the highly supersonic regime. It can also be shown that for the weakly compressible or incompressible regime, it is the second term that dominates.
@@
For the application to interstellar turbulence, we consider highly supersonic turbulence. We first verify the assumptions in our formulation using data from a simulation of supersonic MHD turbulence with a sonic Mach number of $10$ (see Section 1.3). The left panel of Fig. 4 confirms that the density-velocity mixed correlation functions are approximately equal to the density correlation function times $v^{\prime 2}$, $B^{\rho,v}_{ll}(s)\approx B^{\rho,v}_{nn}(s)\approx B_{\rho}(s)v^{\prime 2}$, as expected from the statistical independence of density and velocity fields. The right panel shows that the density-weighted velocity structure functions exhibit similar behaviors as the velocity structure functions, $S^{\rm dw}_{ij}(s)\approx S_{ij}(s)$, except for a slightly larger normalization of the transversal structure function, $S_{nn}$ (filled blue squares), and a slightly steeper slope of the density-weighted transversal structure function, $S^{\rm dw}_{nn}$ (empty red squares). In addition, the left panel shows that the density correlation function (black circles) can be approximated by a power-law function. A least-squares fit for the density correlation function, $B_{\rho}(s)\propto s^{-\beta}$, gives $\beta=0.61$, which is the value we adopt in our applications. Assuming that $B_{\rho}(s)=cs^{-\beta}$, we may integrate Equation (16) to obtain,
@@
which is found to be in excellent agreement with results from our simulation of highly supersonic turbulence (see Section 1.3). Figure 5 shows the scaling of $\langle\bm{j}^{2}\rangle^{1/2}$ from both the simulation (blue squares) and the analytical result (red dashed line). Both the slope and the normalization are nearly identical in the two cases, for distances within the limited inertial range of the turbulence in the simulation. Small deviations appear only at large scale, affected by the driving force, and small scale, affected by numerical dissipation.
Figure 5: Scaling of the variance of the specific angular momentum, $\langle j^{2}\rangle^{1/2}$, as a function of the size $R$ in units of half the box size, $L$. The blue square symbols and the connected blue lines are the results averaged over four snapshots of the simulation. The red dashed line is the linear prediction from Equation 21.
We offer a more intuitive derivation of the linear scaling of $\langle\bm{j}^{2}\rangle^{1/2}$ in highly supersonic turbulence, Equation 21, where the role of the mass center offset is more easily appreciated. We consider $R$ much smaller than the integral scale of turbulence, so that one may neglect fluctuations of ${\bm{v}}$ and assume it is constant within the sphere, ${\bm{v}}={\bm{v}}_{\rm c}$. Setting ${\bm{v}}={\bm{v}}_{\rm c}$ in Equation (11) for $\bm{J}$ yields
@@
where $M$ is the total mass $M$ (Equation 10) and $\bm{r}_{c}$ is the mass center defined as $\bm{r}_{c}=\frac{1}{M}\int\rho\bm{r}\exp\left(-\frac{r^{2}}{R^{2}}\right)d^{3}r$. Due to strong density fluctuations in the highly supersonic turbulence, $\bm{r}_{c}$ may deviate significantly from the geometric center of the sphere, leading to considerable angular momentum, as implied by equation (22). The variance of $\bm{r}_{c}$ may be estimated as,
@@ 1.2 Angular Momentum Versus Size for Molecular Clouds
In the observational studies, the specific angular momentum of MCs is computed as $j=R^{2}\Omega$, where $\Omega$ is the cloud’s overall angular speed, derived from the gradient of the mean radial velocity of emission line spectra at different cloud positions. Here we show that this definition of $j$, is equivalent to our definition in Equation 11, if the gas density is assumed to be constant, hence the observed scaling of $j$ in MCs can be predicted from the $\langle{\bm{J}}^{2}\rangle_{2}$ term in our formalism in Section 1.1, imposing constant density.
@@
which is equivalent to Equation 11 assuming constant density. The observational method is thus equivalent to setting the density to a constant value in our formalism in Section 1.1. In that case, the density correlation function is $B_{\rho}\simeq\rho_{0}^{2}$, where $\rho_{0}$ is the constant density of the flow, hence $\beta=0$. As a result, the first of the two contributions to the variance of the angular momentum in this case is $\langle\bm{J}^{2}\rangle_{1}=0$. In addition, the second term, $\langle\bm{J}^{2}\rangle_{2}$, is simplified by setting $\beta=0$ in Equation 20. To evaluate the velocity structure functions, we make use of Larson’s velocity-size relation [54], $\Delta v(\ell)=C\ell^{\alpha}$, where $\alpha\simeq 0.5$ [42] and $C$ is a constant, and assume equipartition between solenoidal and compressive modes in the power spectrum, so that $S_{ll}(s)=S_{nn}(s)=\frac{1}{3}C^{2}s^{2\alpha}$. Thus, Equation 20 is further simplified by setting $c_{l}=c_{n}=C^{2}/3$, and we obtain:
@@
If $\alpha=0.5$, we have $\langle\bm{j}^{2}\rangle^{1/2}=(18\pi)^{-\frac{1}{4}}CR^{\frac{3}{2}}$. Using the relation from Solomon et al. [42], $\Delta v(\ell)=0.72(\ell/1$ pc $)^{0.5}$ km s-1 (where $\Delta v(\ell)$ is an estimate of the three-dimensional velocity dispersion), we find that
@@
As shown in Figure 1 in the Main text, this predicted scaling is almost identical to the relation between specific angular momentum and size of MCs derived from the observational data. The scaling of the angular momentum of MCs and dense cores is not directly applicable to our problem of estimating the angular momentum of the gas captured by a PMS star along its trajectory. If applied to our problem, it would significantly underestimate the specific angular momentum of the gas captured by the star, as shown by Figure 1.
@@ 1.3 Numerical Simulation and $j$ Scaling
In order to test the analytical results, we use a numerical simulation of randomly-driven, supersonic, magneto-hydrodynamic (MHD) turbulence, designed to simulate the star-formation process in a turbulent interstellar cloud. The simulation is the same used in Kuffmeier et al. [55], also equivalent to the high reference simulation in Haugbølle et al. [56], except that the numerical resolution (the root grid) is larger by a factor of two. The reader is referred to Haugbølle et al. [56] for details of the numerical methods, which is only briefly summarized here. The simulation solves the MHD equations with the adaptive-mesh-refinement (AMR) code Ramses [57], with a root grid of $512^{3}$ cells and six levels of refinement, corresponding to a smallest cell of size $\Delta x=25$ AU for the assumed box size of 4 pc. The total mass, mean density, and mean magnetic field strength are $M_{\rm box}=3000\rm\,M_{\odot}$, ${\bar{n}_{\rm H}}=1897\,\rm cm^{-3}$, and ${\bar{B}}=7.2\,\mu\rm G$, appropriate for a typical star-forming cloud at that scale. The equation of state is assumed to be isothermal, and the boundary conditions are periodic. The turbulence is first driven, without self-gravity, for $\sim 20$ dynamical times, with a random solenoidal acceleration giving an rms sonic Mach number of approximately 10. The simulation is then continued for $\sim 2$ Myr with self-gravity and sink particles to capture the formation of individual stars, yielding 317 stars with a mass distribution consistent with the observed stellar IMF [58, 59].
@@
To test Equation 21, we generate density and velocity snapshots in a uniform grid of $512^{3}$ cells (using only the root grid of the AMR simulation), and compute the angular momentum, ${\bm{J}}$, and the mass, $M$, within spherical volumes of radius $R$ with a Gaussian cutoff, consistent with Equations 11 and 10. We then compute the specific angular momentum, ${\bm{j}}={\bm{J}}/M$, for each sphere. We use $16^{3}$ spheres with centers uniformly distributed in the computational domain, and collect data from six snapshots, so the variance of ${\bm{j}}$ is computed from an average over 16,384 spherical volumes. The procedure is repeated for 7 values of the cutoff radius, $R=4$, 8, 16, 32, 64, 128, 256×$\times\Delta x$. The result is shown by the filled blue squares in Figure 5. The least-squares fit (solid black line in Figure 5) has both slope and normalization indistinguishable from those predicted by Equation 21 (red dashed line in Figure 5).
@@
The same 6 snapshots were also used to test the key assumption of the derivation of Equation 21, that is the independence of density and velocity fields, and to measure the scaling exponent ($\beta=0.61$) of the density correlation function, $B_{\rho}(s)$ (see Section 1.1).

### v2

@@ 1.1 Angular-Momentum Scaling in a Turbulent Flow
where the subscripts “1” and “2” indicate quantities at two points $\bm{r}_{1}$ and $\bm{r}_{2}$, respectively. Under the assumption of statistical homogeneity and isotropy, it is straightforward to show that $\langle\rho_{1}\rho_{2}v_{1i}v_{2j}\rangle=B^{\rho,v}_{ij}(\bm{s})-\frac{1}{2}B_{\rho}S^{\rm dw}_{ij}(\bm{s})$, where the density correlation function $B_{\rho}(\bm{s})\equiv\langle\rho_{1}\rho_{2}\rangle$, the mixed correlation function $B^{\rho,v}_{ij}(\bm{s})\equiv\langle\rho_{1}\rho_{2}v_{1i}v_{1j}\rangle$ and the density-weighted structure function $S^{\rm dw}_{ij}(\bm{s})\equiv\langle\rho_{1}\rho_{2}(v_{2i}-v_{1i})(v_{2j}-v_{1j})\rangle/B_{\rho}$, depend on the separation, $\bm{s}=\bm{r}_{2}-\bm{r}_{1}$. Due to the nonlinear and stochastic nature of turbulence, the relative phase of the velocity and density fields appears to be random, leading to a negligible density-velocity correlation in highly supersonic turbulence, as confirmed by our simulation data (shown below) and other numerical studies [e.g. 51]. If the density and velocity fields are assumed to be independent, we have $B^{\rho,v}_{ij}=\langle\rho_{1}\rho_{2}\rangle\langle v_{1i}v_{1j}\rangle=B_{\rho}v^{\prime 2}\delta_{ij}$ with $v^{\prime}$ the 1-dimensional rms velocity, suggesting that the longitudinal and transverse correlation functions can be written as $B^{\rho,v}_{ll}=B^{\rho,v}_{nn}=B_{\rho}v^{\prime 2}$ (see Supplementary Figure 1 and discussion below). The above equation can then be rewritten as,
@@
The two contributions in Equation (13) can be intuitively understood as follows. The first term which arises mainly due to density fluctuations represents the offset of the mass center from the geometric center of sphere (see below), while the second term, which depends on the velocity structure function, originates from the “imbalance” of the turbulent velocity on the opposite sides of the geometric center, leading to a “residual” angular momentum. For convenience, we denote the two terms as $\langle\bm{J}^{2}\rangle_{1}$ and $\langle\bm{J}^{2}\rangle_{2}$, respectively. As discussed below, the first term provides the dominant contribution in the highly supersonic regime. It can also be shown that for the weakly compressible or incompressible regime, it is the second term that dominates.
@@
For the application to interstellar turbulence, we consider highly supersonic turbulence. We first verify the assumptions in our formulation using data from a simulation of supersonic MHD turbulence with a sonic Mach number of $10$ (see Section 1.3). The left panel of Supplementary Figure 1 confirms that the density-velocity mixed correlation functions are approximately equal to the density correlation function times $v^{\prime 2}$, $B^{\rho,v}_{ll}(s)\approx B^{\rho,v}_{nn}(s)\approx B_{\rho}(s)v^{\prime 2}$, as expected from the statistical independence of density and velocity fields, as mentioned above. The right panel shows that the density-weighted velocity structure functions exhibit similar behaviors as the velocity structure functions, $S^{\rm dw}_{ij}(s)\approx S_{ij}(s)$, except for a slightly larger amplitude of the transversal structure function, $S_{nn}$ (filled blue squares), and a slightly steeper slope of the density-weighted transversal structure function, $S^{\rm dw}_{nn}$ (empty red squares). In addition, the left panel shows that the density correlation function (black circles) can be approximated by a power-law function. A least-squares fit for the density correlation function, $B_{\rho}(s)\propto s^{-\beta}$, gives $\beta=0.61$, which is the value we adopt in our applications. Assuming that $B_{\rho}(s)=cs^{-\beta}$, we may integrate Equation (16) to obtain,
@@
which is found to be in excellent agreement with results from our simulation of highly supersonic turbulence (see Section 1.3). Supplementary Figure 2 shows the scaling of $\langle\bm{j}^{2}\rangle^{1/2}$ from both the simulation (blue squares) and the analytical result (red dashed line). Both the slope and the normalization are nearly identical in the two cases, for distances within the limited inertial range of the turbulence in the simulation. Small deviations appear only at large scale, affected by the driving force, and small scale, affected by numerical dissipation.
We offer a more intuitive derivation of the linear scaling of $\langle\bm{j}^{2}\rangle^{1/2}$ in highly supersonic turbulence, Equation (21), where the role of the mass center offset is more easily appreciated. We consider $R$ much smaller than the integral scale of turbulence, so that one may neglect fluctuations of ${\bm{v}}$ and assume it is constant within the sphere, ${\bm{v}}={\bm{v}}_{\rm c}$. Setting ${\bm{v}}={\bm{v}}_{\rm c}$ in Equation (11) for $\bm{J}$ yields
@@
where $M$ is the total mass $M$ (Equation (10)) and $\bm{r}_{c}$ is the mass center defined as $\bm{r}_{c}=\frac{1}{M}\int\rho\bm{r}\exp\left(-\frac{r^{2}}{R^{2}}\right)d^{3}r$. Due to strong density fluctuations in the highly supersonic turbulence, $\bm{r}_{c}$ may deviate significantly from the geometric center of the sphere, leading to considerable angular momentum, as implied by Equation (22). The variance of $\bm{r}_{c}$ may be estimated as,
@@ 1.2 Angular Momentum Versus Size for Molecular Clouds
In the observational studies, the specific angular momentum of MCs is computed as $j=R^{2}\Omega$, where $\Omega$ is the cloud’s overall angular speed, derived from the gradient of the mean radial velocity of emission line spectra at different cloud positions. Here we show that this definition of $j$ is equivalent to our definition in Equation (11), if the gas density is assumed to be constant, hence the observed scaling of $j$ in MCs can be predicted from the $\langle{\bm{J}}^{2}\rangle_{2}$ term in our formalism in Section 1.1, imposing constant density.
@@
which is equivalent to Equation (11) assuming constant density. The observational method is therefore equivalent to setting the density at a constant value in our formalism in Section 1.1. In that case, the density correlation function is $B_{\rho}\simeq\rho_{0}^{2}$, where $\rho_{0}$ is the constant density of the flow, hence $\beta=0$. As a result, the first of the two contributions to the variance of the angular momentum in this case is $\langle\bm{J}^{2}\rangle_{1}=0$. In addition, the second term, $\langle\bm{J}^{2}\rangle_{2}$, is simplified by setting $\beta=0$ in Equation (20). To evaluate the velocity structure functions, we make use of Larson’s velocity-size relation [52], $\Delta v(\ell)=C\ell^{\alpha}$, where $\alpha\simeq 0.5$ [53] and $C$ is a constant, and assume equipartition between solenoidal and compressive modes in the power spectrum, so that $S_{ll}(s)=S_{nn}(s)=\frac{1}{3}C^{2}s^{2\alpha}$. Thus, Equation (20) is further simplified by setting $c_{l}=c_{n}=C^{2}/3$, and we obtain:
@@
If $\alpha=0.5$, we have $\langle\bm{j}^{2}\rangle^{1/2}=(18\pi)^{-\frac{1}{4}}CR^{\frac{3}{2}}$. Using the relation from Solomon et al. (1987) [53], $\Delta v(\ell)=0.72(\ell/1$ pc $)^{0.5}$ km s-1 (where $\Delta v(\ell)$ is an estimate of the three-dimensional velocity dispersion), we find that
@@
As shown in the left panel of Figure 1 in the Main text, this predicted scaling is almost identical to the relationship between the specific angular momentum and the size of MCs derived from the observational data. The scaling of the angular momentum of MCs and dense cores is not directly applicable to our problem of estimating the angular momentum of the gas captured by a PMS star along its trajectory. If applied to our problem, it would significantly underestimate the specific angular momentum of the gas captured by the star, as shown by the left panel of Figure 1. The two scaling laws shown in that figure are comparable at the turbulence outer scale (the computational box size in this case), where $\sigma_{v,0}$ (or the corresponding $\sigma_{v,\rm rel}$) is computed, because at that scale the velocity fluctuations (as well as net rotation) are of the order of $\sigma_{v,0}$. With a larger outer scale, the normalization of the linear scaling law would be larger, further increasing the discrepancy with the MC’s scaling law at small scales.
@@ 1.3 Numerical Simulation and $j$ Scaling
In order to test the analytical results, we use a numerical simulation of randomly driven, supersonic, magnetohydrodynamic (MHD) turbulence, designed to simulate the star-formation process in a turbulent interstellar cloud. The simulation, also used in other recent works [54, 55, 37, 56, 44], is equivalent to the high reference simulation in Haugbølle et al. (2018) [57], except that the numerical resolution (the root grid) is larger by a factor of two (a factor of 8 increase in the number of cells). The reader is referred to that work for details of the numerical methods, which is only briefly summarized here. The simulation solves the MHD equations with the adaptive mesh-refinement (AMR) code Ramses [58], with a root grid of $512^{3}$ cells and six levels of refinement, corresponding to a smallest cell of size $\Delta x=25$ AU for the assumed box size of 4 pc. The total mass, mean density, and mean magnetic field strength are $M_{\rm box}=3000\rm\,M_{\odot}$, ${\bar{n}_{\rm H}}=1897\,\rm cm^{-3}$, and ${\bar{B}}=7.2\,\mu\rm G$, appropriate for a typical star-forming cloud at that scale. The equation of state is assumed to be isothermal, and the boundary conditions are periodic. The turbulence is first driven, without self-gravity, for $\sim 20$ dynamical times, with a random solenoidal acceleration giving an rms sonic Mach number of approximately 10. Keeping the driving force, the simulation is then continued for $\sim 2$ Myr with self-gravity and sink particles to capture the formation of individual stars, yielding 317 stars with a mass distribution consistent with the observed stellar IMF, and a final star formation efficiency of approximately 8%. [59, 60]. Feedback from high-mass stars is not included, as only two stars with mass close to $10\,M_{\odot}$ are found towards the end of the simulation, and their structure should probably be inflated due to the high accretion rate, thus preventing a significant UV flux at this early stage. In addition, all the nearby star-forming regions whose disk sizes are reported here are also devoid of high-mass stars.
@@
To test Equation (21), we generate density and velocity snapshots in a uniform grid of $512^{3}$ cells (using only the root grid of the AMR simulation), and compute the angular momentum, ${\bm{J}}$, and the mass, $M$, within spherical volumes of radius $R$ with a Gaussian cutoff, consistent with Equations (11) and (10). We then compute the specific angular momentum, ${\bm{j}}={\bm{J}}/M$, for each sphere. We use $16^{3}$ spheres with centers uniformly distributed in the computational domain, and collect data from six snapshots, so the variance of ${\bm{j}}$ is computed from an average over 16,384 spherical volumes. The procedure is repeated for 7 values of the cutoff radius, $R=4$, 8, 16, 32, 64, 128, 256×$\times\Delta x$. The result is shown by the filled blue squares in Supplementary Figure 2. The least-squares fit (solid black line in Figure 4) has both slope and normalization indistinguishable from those predicted by Equation (21) (red dashed line in Figure 4).
@@
The same 6 snapshots were also used to test the key assumption of the derivation of Equation (21), that is the independence of density and velocity fields, and to measure the scaling exponent ($\beta=0.61$) of the density correlation function, $B_{\rho}(s)$ (see Section 1.1).
@@ 1.4 Disk Model
1.5 Observed Disk Sizes
The stellar mass and disk size values (and the corresponding disk $j$ values) shown in the two plots of Figure 1 are all measured from PMS stars (i.e. Class II sources). They are taken from four different publications [33, 42, 35, 43], avoiding duplication when the same objects appear in different samples. For the sake of statistics, the disk radii are measured from dust continuum observations, $R_{\rm mm}$, rather than gas observations, $R_{\rm CO}$, as the latter are available only for a smaller number of disks. The values of $R_{\rm mm}$ are taken from the published tables, and are in all cases defined as the radius containing 90% of the sub-millimeter flux of the disk. It is well established that PDs appear to be a few times larger when traced by gas rather than dust, which is often interpreted as evidence of growth and radial settling of dust grains. For example, for the disks from Long et al. (2022) [35] that have both types of sizes, $R_{\rm CO}/R_{\rm mm}=2.9\pm 1.2$, on average. However, the ratio does not depend on stellar mass, hence adopting $R_{\rm CO}$ instead of $R_{\rm mm}$ in our work would not change our conclusions on the mass dependence. The derived $j$ depends on the $R_{\rm d}^{1/2}$, as shown by Equation (29), so it would increase on average by only 70%, without affecting our conclusion that BH accretion provides a large enough angular momentum to explain the observed disk sizes. To illustrate the comparison of our analytical and numerical results with the observations based only on the gas radii, $R_{\rm CO}$, we have made a second version of the right panel of Figure 1, where the values of $j$ for the disks are based on their $R_{\rm CO}$ (Supplementary Figure 3). The dependence of the disk $j$ on the stellar mass is essentially the same as for $R_{\rm mm}$ (the slope of the least-squares fit is the same within the $1\sigma$ uncertainty), and the $j$ values are mostly within the analytical prediction for ages between 1 and 4 Myr, with their line of best fit very close to that for the numerical data points.